I've decided that the mentor group should be made up of 3 girls and 3 boys. Of the students who have A's 9 of them are girls and 6 are boys. How many ways can I form a group of 3 boys and 3 girls?









Answer by stanbon(75887)

the mentor group should be made up of 3 girls and 3 boys. Of the students who have A's 9 of them are girls and 6 are boys. How many ways can I form a group of 3 boys and 3 girls?
---------------
Ans:
# of ways to pick the 3 girls: 9C3
# of ways to pick the 3 boys: 6C3
---------
Total # of 9C3 * 6C3 = 84*20 = 1680
==============
Cheers,
Stan H.